# A constructive approach to strengthen algebraic descriptions of function and operator classes

### v6

@@ 1 Introduction
Another description of $\mathcal{F}_{{0,L}}(\mathbb{R}^{d})$ consists in juxtaposing convexity and another definition of smoothness, obtained by integration of the Lipschitz continuity of the gradients, see, e.g., [nesterov2018lectures, Theorem 2.1.5], and imposing, $\forall x,y\in\mathbb{R}^{d}$,
@@ 1.1 Related works
Positive polynomials: SOS and co-positive formulations.
@@
Setting $\gamma=\theta^{2}$, enforcing $P(\theta)\leqslant 0,\ \forall\theta\geqslant 0$ is equivalent to enforcing a $n$-dimensional polynomial of degree $2k$, $\bar{P}(\gamma)\in\mathbb{R}_{2k}[\gamma_{1},\ldots,\gamma_{n}]$ to be non-positive everywhere. Interest in non-negative polynomials can be traced back to 1888 [hilbert1888darstellung], and they have been widely studied since; see, e.g., [lasserre2009moments, marshall2008positive] for surveys. While the set of non-negative polynomials is non-tractable in general, it can be relaxed in several tractable ways, e.g., by requiring polynomials to be sum of squares (SOS). That is, there exists a matrix $M\succcurlyeq 0$ such that $\bar{P}(\gamma)=-X^{T}MX$, where $X$ is a vector of size $\binom{k+n}{n}$ containing all monomials of degree $\leqslant k$ in $\gamma$. It is well-known that for the special cases $(n=1,k\in\mathbb{N})$, quadratic polynomials and $(n=2,k=2)$, non-negative and SOS polynomials are equivalent [hilbert1888darstellung]. In any case, the set of SOS polynomials provides a natural path to tractable convex relaxations of positive polynomials in general.
Alternatively, in some cases, enforcing $P(\theta)\leqslant 0,\ \forall\theta\geqslant 0$ can be equivalently formulated as requiring existence of a co-positive matrix $M$ (i.e., $x^{T}Mx\geqslant 0,\ \forall x\geqslant 0$), such that $P(\theta)=-X(\theta)^{T}MX(\theta)$ for some vector $X(\theta)$. Co-positive programming has been widely investigated since 1952 [motzkin1952copositive], see, e.g., [dur2010copositive] for a survey. While certifying the co-positivity of a matrix is generally $NP$-hard, it is semi-definite representable for small problem dimensions ($\leqslant 4$, see, e.g., [maxfield1962matrix]) and can naturally be relaxed to tractable convex representations otherwise.
@@ 1.2 Contributions
The contribution of this work is twofold. First, Section 2 presents the refinement technique, and illustrates it on the class of smooth convex functions, providing an original approach (not relying on integration arguments) to obtain $p^{\prime\prime}_{0,L}$ from $p_{0,L}$. Then, Section 3 exploits the technique to obtain new conditions for function and operator classes that have sparked interest in the optimization community (see, e.g., [abbaszadehpeivasti2022conditionslinearconvergencegradient, bolte2010characterizations, drori2018properties, karimi2016linear, nesterov2012efficiency, nesterov2015universal, nesterov2018lectures, ryu2020operator]) and for which no extension condition yet exists. These refined conditions are obtained via a single round of the refinement procedure, and are hence a priori only necessary extension conditions. However, we show on several examples how they quantitatively improve upon classical descriptions. In addition, we provide, for most cases, SDP representations (in terms of the conditions imposed on a dataset) of those refined conditions. In particular, we strengthen existing descriptions of
@@
  • convex blockwise smooth functions, see, e.g., [beck, nesterov2012efficiency] for motivations, and [abbaszadehpeivasti2022convergence, kamri2023worst] for recent analysis of the cyclic or randomized blockwise coordinate descent methods on this class: SDP formulation (Section 3.2),
  • Lipschitz strongly monotone operators, see, e.g., [bauschke2017convex, ryu2020operator] for motivations, and [gorbunov2022extragradient, gorbunov2022last, korpelevich1976extragradient] for recent analysis of the extra gradient and optimistic gradient on this class: SDP formulation (Section 3.3),
@@ 2.2 One-point strengthenings
Consider a property $p:(\mathbb{R}^{d}\times\mathbb{R}^{D})^{m}\to\mathbb{R}$, $\bar{Q}\subseteq\mathbb{R}^{d}$, and $K\in\mathbb{N},m\leqslant K\leqslant|\bar{Q}|$. We define $\tilde{p}_{K,\bar{Q}}:(\mathbb{R}^{d}\times\mathbb{R}^{D})^{K}\to\mathbb{R}$, the $K$-wise one-point strengthening of $p$ with respect to $\bar{Q}$, as
@@ 3.1 Extension of smooth functions satisfying a Łojasiewicz condition
(6) $M^{ij}=\begin{bmatrix}-2\mu A&0&M_{13}&M_{14}\\ 0&M_{22}&-M_{14}&M_{24}\\ M_{13}&-M_{14}&M_{33}&0\\ M_{14}&M_{24}&0&-2\mu A-\frac{2\mu(B-C-(L+3\mu)A)}{2L+\mu}-\frac{4\mu^{2}((L+\mu)A-2B)}{(2L+\mu)^{2}}-\frac{8\mu^{3}B}{(2L+\mu)^{3}}\end{bmatrix}\succcurlyeq 0,$
@@
The detailed proof of Proposition 5 is deferred to Appendix B.2, and can be summarized as follows. The last condition in (5) can be formulated as a one-dimensional polynomial of degree 33 $P(\alpha)\in\mathbb{R}_{3}[\alpha]$, whose coefficients depend on $\{(x_{i},f_{i},g_{i})\}_{i=1,2,\star}$, which must be non-positive for all $\alpha\in[0,\frac{2\mu}{2L+\mu}]$. Setting $\alpha=\frac{2\mu}{2L+\mu}\frac{\theta^{2}}{1+\theta^{2}}$, this can be equivalently expressed as requiring a one dimensional polynomial of degree 66 $\bar{P}(\theta)\in\mathbb{R}_{6}[\theta]$ to be non-positive for all $\theta\in\mathbb{R}$. Since $\bar{P}(\theta)$ is one-dimensional, requiring its non-positivity is strictly equivalent to requiring $\bar{P}(\theta)$ to be sum of squares [hilbert1888darstellung], which can be expressed as an SDP.
@@ 3.2 Convex blockwise smooth functions
(8) $M_{m}^{ijk}=\begin{bmatrix}-A_{ikm}&0&M_{13}\\ 0&M_{22}&0\\ M_{13}&0&-A_{ijm}\end{bmatrix}\succcurlyeq 0,$
@@
The detailed proof of Proposition 7 is deferred to Appendix C.2, and can be summarized as follows. Condition (7) is a one-dimensional polynomial of degree $2$, $P(\lambda)\in\mathbb{R}_{2}[\lambda]$, whose coefficients depend on $\{(x_{i},f_{i},g_{i})\}_{i=1,2,3}$, which must be non-positive for all $\lambda\in[0,1]$. Setting $\lambda=\frac{\theta^{2}}{1+\theta^{2}}$, this can be equivalently expressed as requiring a one-dimensional polynomial of degree $4$, $\bar{P}(\theta)\in\mathbb{R}_{4}[\theta]$ to be non-positive for all $\theta\in\mathbb{R}$. Since $\bar{P}(\theta)$ is one-dimensional, requiring its non-positivity is strictly equivalent to requiring $\bar{P}(\theta)$ to be a sum of squares [hilbert1888darstellung], which can be expressed as an SDP.
@@ 3.3 Strongly monotone Lipschitz operators
With the same notation as in Proposition 9 and considering a triplet $S=\{(x_{i},t_{i})\}_{i\in\llbracket 3\rrbracket}$, it holds that $S$ satisfies (10) only if  $\forall i,j,k\in\llbracket 3\rrbracket$, there exists some $M_{A,B}^{ijk}\succcurlyeq 0$, $M_{B,A}^{ijk}\succcurlyeq 0$, where
(11) $M_{A,B}^{ijk}=\begin{bmatrix}-B_{ij}&0&0&M_{14}&M_{15}&M_{16}&M_{17}\\ 0&-A_{jk}&-M_{15}&0&-M_{14}&M_{26}&M_{27}\\ 0&-M_{15}&-B_{ij}&M_{34}&-M_{16}&0&M_{37}\\ M_{14}&0&M_{34}&-B_{jk}&-M_{27}&M_{46}&0\\ M_{15}&-M_{14}&-M_{16}&-M_{27}&M_{55}&-M_{37}&-M_{46}\\ M_{16}&M_{26}&0&M_{46}&-M_{37}&-A_{ik}&0\\ M_{17}&M_{27}&M_{37}&0&-M_{46}&0&-B_{ik}\end{bmatrix}$,
@@
$M_{55}+2M_{17}+2M_{26}+2M_{34}$ $=A_{ij}+2\mu B_{ij}-A_{jk}-A_{ik}.$
The detailed proof of Proposition 10 is deferred to Appendix D.2, and can be summarized as follows. Condition (10) is a 33dimensional polynomial of degree $3$, $P(\lambda_{k},\mu_{i},\mu_{j})\in\mathbb{R}_{3}[\lambda_{k},\mu_{i},\mu_{j}]$, whose coefficients depend on $\{(x_{i},t_{i})\}_{i=1,2,3}$, which must be non-positive for all $\lambda_{k},\mu_{i},\mu_{j}\geqslant 0$. Setting $\lambda_{k}=\theta^{2},\mu_{i}=\alpha^{2},\mu_{j}=\gamma^{2}$, this can be equivalently expressed as requiring a $3$-dimensional polynomial of degree $6$, $\bar{P}(\theta,\alpha,\gamma)\in\mathbb{R}_{6}[\theta,\alpha,\gamma]$ to be non-positive for all $\theta,\alpha,\gamma\in\mathbb{R}$. Requiring non-positivity of $\bar{P}(\theta,\alpha,\gamma)$ can be relaxed as requiring $\bar{P}(\theta,\alpha,\gamma)$ to be sum of squares [hilbert1888darstellung], which can be expressed as an SDP.
@@ 4 Concluding remarks
As this work is strongly motivated by algorithm analyses—it was recently shown that interpolation/extension conditions are key for systematically analyzing their performance; see, e.g., [taylor2017smooth] or recent presentations [goujaud2023fundamental, rubbens2023interpolation]—, we provide a few examples of such applications using the conditions obtained in Section 3. These include analysis of the gradient method on smooth function satisfying a Łojasiewisz condition [bolte2017error, abbaszadehpeivasti2022conditionslinearconvergencegradient, karimi2016linear] (see Section 3.1), of the random or cyclic coordinate descent on convex blockwise smooth functions [abbaszadehpeivasti2022convergence, beck, kamri2023worst, nesterov2012efficiency] (see Section 3.2), and of the optimistic gradient and extragradient on Lipschitz monotone operators [gorbunov2022last, gorbunov2022extragradient, korpelevich1976extragradient] (see Section 3.3).
@@ B.2 Derivation of a SOS reformulation: proof of Proposition 5
$\Leftrightarrow$ $0\geqslant a_{0}+a_{1}\alpha+a_{2}\alpha^{2}+a_{3}\alpha^{3},\ \forall\alpha\in[0,\frac{2\mu}{2L+\mu}],$
@@
Setting $\alpha=\frac{2\mu}{2L+\mu}\frac{\theta^{2}}{1+\theta^{2}}$, and multiplying by $(1+\theta^{2})^{3}$ yields
$0\geqslant a_{0}(1+\theta^{2})^{3}+a_{1}\frac{2\mu}{2L+\mu}\theta^{2}(1+\theta^{2})^{2}+a_{2}\frac{4\mu^{2}}{(2L+\mu)^{2}}\theta^{4}(1+\theta^{2})+a_{3}\frac{8\mu^{3}}{(2L+\mu)^{3}}\theta^{6}:=P(\theta),\ \forall\theta\in\mathbb{R},$
which can be checked as a semi-definite program. Indeed, $P(\theta)$ is a polynomial of one variable $\theta$, which is negative for all $\theta\in\mathbb{R}$ if and only if there exists $M\preccurlyeq 0$ such that
$P(\theta)=\begin{pmatrix}1\\ \theta\\ \theta^{2}\\ \theta^{3}\end{pmatrix}^{\top}M\begin{pmatrix}1\\ \theta\\ \theta^{2}\\ \theta^{3}\end{pmatrix}$,
which can be stated linearly in terms of the coefficients of $M$ and those of $P(\cdot)$. Denoting the $i,j^{\text{th}}$ entrance of $M$ by $M_{i,j}$, $i,j\in\llbracket 4\rrbracket$, the list of constraints is given by
$M_{1,1}$ $=a_{0}$
$M_{2,2}+2M_{1,3}$ $=3a_{0}+a_{1}\frac{2\mu}{2L+\mu}$
$M_{3,3}+2M_{2,4}$ $=3a_{0}+2a_{1}\frac{2\mu}{2L+\mu}+a_{2}\frac{4\mu^{2}}{(2L+\mu)^{2}}$
$M_{4,4}$ $=a_{0}+a_{1}\frac{2\mu}{2L+\mu}+a_{2}\frac{4\mu^{2}}{(2L+\mu)^{2}}+a_{3}\frac{8\mu^{3}}{(2L+\mu)^{3}}$
$M_{1,2}=M_{3,4}=M_{1,4}+M_{2,3}$ $=0$
@@ C.2 Derivation of a SOS reformulation: proof of Proposition 7
$0\geqslant$ $(1-\lambda)A_{ijm}+\lambda A_{ikm}+\lambda(1-\lambda)B_{jkm},\ \forall m\in\llbracket n\rrbracket,\ \forall\lambda\in[0,1],$
@@
Setting $\lambda=\frac{\theta^{2}}{1+\theta^{2}}$, and multiplying by $(1+\theta^{2})^{2}$ yields
$0\geqslant$ $(1+\theta^{2})A_{ijm}+\theta^{2}(1+\theta^{2})A_{ikm}+\theta^{2}B_{jkm}:=P_{m}(\theta),\ \forall m\in\llbracket n\rrbracket,\ \forall\lambda\in[0,1],$
which can again be checked as a semi-definite program. Indeed, $P_{m}(\theta)$ is a polynomial of one variable $\theta$, which is negative for all $\theta\in\mathbb{R}$ if and only if there exists $M_{m}\preccurlyeq 0$ such that
$P_{m}(\theta)=\begin{pmatrix}1\\ \theta\\ \theta^{2}\end{pmatrix}^{\top}M_{m}\begin{pmatrix}1\\ \theta\\ \theta^{2}\end{pmatrix}$,
which can be stated linearly in terms of the coefficients of $M_{m}$ and those of $P_{m}(\cdot)$. Denoting the $i,j^{\text{th}}$ entrance of $M$ by $M_{i,j}$, $i,j\in\llbracket 3\rrbracket$, the list of constraints is given by
$M_{1,1}$ $=A_{ikm}$
$M_{2,2}+2M_{1,3}$ $=A_{ikm}+A_{ijm}+B_{jkm}$
$M_{3,3}$ $=A_{ijm}$
$M_{1,2}=M_{2,3}$ $=0.$

### v7

@@ 1 Introduction
Another description of $\mathcal{F}_{{0,L}}(\mathbb{R}^{d})$ consists in juxtaposing convexity and another possible definition of smoothness, obtained by integration of the Lipschitz continuity of the gradients, see, e.g., [nesterov2018lectures, Theorem 2.1.5], and imposing, $\forall x,y\in\mathbb{R}^{d}$,
@@ 1.1 Related works
Positive polynomials: sum of squares (SOS) and co-positive formulations.
@@
Setting $\gamma=\theta^{2}$, enforcing $P(\theta)\leqslant 0,\ \forall\theta\geqslant 0$ is equivalent to enforcing a $n$-dimensional polynomial of degree $2k$, $\bar{P}(\gamma)\in\mathbb{R}_{2k}[\gamma_{1},\ldots,\gamma_{n}]$ to be non-positive everywhere. Interest in non-negative polynomials can be traced back to 1888 [hilbert1888darstellung], and they have been widely studied since; see, e.g., [lasserre2009moments, marshall2008positive] for surveys. While the set of non-negative polynomials is non-tractable in general, it can be relaxed in several tractable ways, e.g., by requiring polynomials to be sum of squares (SOS). That is, there exists a matrix $M\succcurlyeq 0$ such that $\bar{P}(\gamma)=-X^{T}MX$, where $X$ is a vector of size $\binom{k+n}{n}$ containing all monomials of degree $\leqslant k$ in $\gamma$. It is well-known that for the special cases $(n=1,k\in\mathbb{N})$, quadratic polynomials and $(n=2,k=2)$, non-negative and SOS polynomials are equivalent [hilbert1888darstellung]. In any case, the set of SOS polynomials provides a natural path to tractable convex relaxations of positive polynomials in general. Alternatively, when $n=1$, one can use the Markov-Lukasc theorem, see, e.g., [szeg1939orthogonal, Theorem 1.21], to express non-negativity of polynomials as SDPs.
Finally, in some cases, enforcing $P(\theta)\leqslant 0,\ \forall\theta\geqslant 0$ can be equivalently formulated as requiring existence of a co-positive matrix $M$ (i.e., $x^{T}Mx\geqslant 0,\ \forall x\geqslant 0$), such that $P(\theta)=-X(\theta)^{T}MX(\theta)$ for some vector $X(\theta)$. Co-positive programming has been widely investigated since 1952 [motzkin1952copositive], see, e.g., [dur2010copositive] for a survey. While certifying the co-positivity of a matrix is generally $NP$-hard, it is semi-definite representable for small problem dimensions ($\leqslant 4$, see, e.g., [maxfield1962matrix]) and can naturally be relaxed to tractable convex representations otherwise.
@@ 1.2 Contributions
The contribution of this work is twofold. First, Section 2 presents the refinement technique, and illustrates it on the class of smooth convex functions, providing an original approach (not relying on integration arguments) to obtain $p^{\prime\prime}_{0,L}$ from $p_{0,L}$. Then, Section 3 exploits the technique to obtain new conditions for function and operator classes that have sparked interest in the optimization community (see, e.g., [abbaszadehpeivasti2022conditionslinearconvergencegradient, bolte2010characterizations, drori2018properties, karimi2016linear, nesterov2012efficiency, nesterov2015universal, nesterov2018lectures, ryu2020operator]) and for which no extension condition yet exists. These refined conditions are obtained via a single round of the refinement procedure, and are hence a priori only necessary extension conditions. However, we show on several examples how they quantitatively improve upon classical descriptions. In addition, we provide, for most cases, semi-definite programs (SDP) representations (in terms of the conditions imposed on a dataset) of those refined conditions. In particular, we strengthen existing descriptions of
@@
  • convex blockwise smooth functions, see, e.g., [beck, nesterov2012efficiency] for motivations, and [abbaszadehpeivasti2022convergence, kamri2023worst] for recent analyses of the cyclic or randomized blockwise coordinate descent methods on this class: SDP formulation (Section 3.2),
  • Lipschitz strongly monotone operators, see, e.g., [bauschke2017convex, ryu2020operator] for motivations, and [gorbunov2022extragradient, gorbunov2022last, korpelevich1976extragradient] for recent analyses of the optimistic and extra gradients on this class: SDP formulation (Section 3.3),
@@ 2.2 One-point strengthenings
Consider a property $p:(\mathbb{R}^{d}\times\mathbb{R}^{D})^{m}\to\mathbb{R}$, $\bar{Q}\subseteq\mathbb{R}^{d}$, and $K\in\mathbb{N},m\leqslant K\leqslant|\bar{Q}|$. We define $\tilde{p}_{K,\bar{Q}}:(\mathbb{R}^{d}\times\mathbb{R}^{D})^{K}\to\mathbb{R}$, the $K$-wise one-point strengthening of $p$, as
@@ 3.1 Extension of smooth functions satisfying a Łojasiewicz condition
(6) $M^{ij}=\begin{bmatrix}M_{11}&M_{12}\\ M_{12}&M_{22}\end{bmatrix}\succcurlyeq 0,\quad\bar{M}^{ij}=\begin{bmatrix}-A(2L+\mu)&\bar{M}_{12}\\ \bar{M}_{12}&\bar{M}_{22}\end{bmatrix}\succcurlyeq 0,$
@@
The detailed proof of Proposition 5 is deferred to Appendix B.2, and relies on Markov-Lukacs theorem, see, e.g., [szeg1939orthogonal, Theorem 1.21].
@@ 3.2 Convex blockwise smooth functions
(8) $M_{m}^{ijk}=\begin{bmatrix}-A_{ijm}&M_{12}\\ M_{12}&M_{22}\end{bmatrix}\succcurlyeq 0,\quad K_{m}^{ijk}\geqslant 0,$
@@
The detailed proof of Proposition 7 is deferred to Appendix C.2, and relies on Markov-Lukacs theorem, see, e.g., [szeg1939orthogonal, Theorem 1.21].
@@ 3.3 Strongly monotone Lipschitz operators
With the same notation as in Proposition 9 and considering a triplet $S=\{(x_{i},t_{i})\}_{i\in\llbracket 3\rrbracket}$, it holds that $S$ satisfies (10) if  $\forall i,j,k\in\llbracket 3\rrbracket$, there exists some $M_{A,B}^{ijk}\succcurlyeq 0$, $M_{B,A}^{ijk}\succcurlyeq 0$, where
(11) $M_{\alpha,\beta}^{ijk}=\begin{bmatrix}-\beta_{ij}&0&0&M_{14}&M_{15}&M_{16}&M_{17}\\ 0&-\alpha_{jk}&-M_{15}&0&-M_{14}&M_{26}&M_{27}\\ 0&-M_{15}&-\beta_{ij}&M_{34}&-M_{16}&0&M_{37}\\ M_{14}&0&M_{34}&-\beta_{jk}&-M_{27}&M_{46}&0\\ M_{15}&-M_{14}&-M_{16}&-M_{27}&M_{55}&-M_{37}&-M_{46}\\ M_{16}&M_{26}&0&M_{46}&-M_{37}&-\alpha_{ik}&0\\ M_{17}&M_{27}&M_{37}&0&-M_{46}&0&-\beta_{ik}\end{bmatrix}$,
@@
$M_{55}+2M_{17}+2M_{26}+2M_{34}$ $=\alpha_{ij}+2\mu\beta_{ij}-\alpha_{jk}-\alpha_{ik}.$
The detailed proof of Proposition 10 is deferred to Appendix D.2.
@@ 4 Concluding remarks
As this work is strongly motivated by algorithm analyses—it was recently shown that interpolation/extension conditions are key for systematically analyzing their performance; see, e.g., [taylor2017smooth] or recent presentations [goujaud2023fundamental, rubbens2023interpolation]—, we provide a few examples of such applications using the conditions obtained in Section 3. These include analysis of the gradient method on smooth function satisfying a Łojasiewisz condition [bolte2017error, abbaszadehpeivasti2022conditionslinearconvergencegradient, karimi2016linear] (see Section 3.1), of the random or cyclic coordinate descent on convex blockwise smooth functions [abbaszadehpeivasti2022convergence, beck, kamri2023worst, nesterov2012efficiency] (see Section 3.2), and of the optimistic gradient and extragradient on Lipschitz monotone operators [gorbunov2022last, gorbunov2022extragradient, korpelevich1976extragradient] (see Section 3.3). In addition, the SDP formulations (6), (8) and (11) of the improved descriptions of (i) smooth functions satisfying a Łojsiewicz condition, (ii) blockwise smooth convex functions, and (iii) (strongly) monotone Lipschitz operators, have been added to the PEPit Python library PEPit documentation [goujaud2024pepit], which allows solving PEPs on these classes straightforwardly.
@@ B.2 Derivation of a SOS reformulation: proof of Proposition 5
$\Leftrightarrow$ $0\geqslant a_{0}+a_{1}\alpha+a_{2}\alpha^{2}+a_{3}\alpha^{3}:=P(\alpha),\ \forall\alpha\in[0,\frac{2\mu}{2L+\mu}],$
@@
By Markov-Lukacs theorem, see, e.g., [szeg1939orthogonal, Theorem 1.21],
$P(\alpha)\leqslant 0,\forall\alpha\in[0,\frac{2\mu}{2L+\mu}]\Leftrightarrow\exists M,\bar{M}\succcurlyeq 0\text{ s.t. }P(\alpha)=-\alpha\begin{bmatrix}1&\alpha\end{bmatrix}M\begin{bmatrix}1\\ \alpha\end{bmatrix}-(\frac{2\mu}{2L+\mu}-\alpha)\begin{bmatrix}1&\alpha\end{bmatrix}\bar{M}\begin{bmatrix}1\\ \alpha\end{bmatrix}$
This condition can be stated linearly in terms of the coefficients of $M,\bar{M}$ and those of $P(\cdot)$. Denoting the $i,j^{\text{th}}$ entrance of $M$ by $M_{ij}$, $i,j\in[2]$, the list of constraints is given by
$\frac{2\mu}{2L+\mu}\bar{M}_{11}$ $=-a_{0},\quad\quad M_{11}-\bar{M}_{11}+\frac{4\mu}{2L+\mu}\bar{M}_{12}=-a_{1}$
$2M_{12}-2\bar{M}_{12}+\frac{2\mu}{2L+\mu}\bar{M}_{22}$ $=-a_{2},\quad\quad\quad\quad\quad\quad\quad\quad\ M_{22}-\bar{M}_{22}=-a_{3}.$
@@ C.2 Derivation of a SOS reformulation: proof of Proposition 7
$0\geqslant$ $(1-\lambda)A_{ijm}+\lambda A_{ikm}+\lambda(1-\lambda)B_{jkm}:=P(\lambda),\ \forall m\in\llbracket n\rrbracket,\ \forall\lambda\in[0,1],$
@@
By Markov-Lukasc theorem [szeg1939orthogonal, Theorem 1.21],
$P(\lambda)\leqslant 0,\forall\lambda\in[0,1]\Leftrightarrow\exists M\succcurlyeq 0,K\geqslant 0\text{ s.t. }P(\lambda)=-\begin{bmatrix}1&\lambda\end{bmatrix}M\begin{bmatrix}1\\ \lambda\end{bmatrix}-\lambda(1-\lambda)K.$
This condition can be stated linearly in terms of the coefficients of $M$, of $K$ and of the coefficients of $P(\cdot)$. Denoting the $i,j^{\text{th}}$ entrance of $M$ by $M_{ij}$, $i,j\in[2]$, the list of constraints is given by
$M_{11}$ $=-A_{ijm},\quad 2M_{12}+K=A_{ijm}-A_{ikm}-B_{jkm},\quad M_{22}-K=B_{jkm}.$